# Concept-Based Interpretable Reinforcement Learning with Limited to No Human Labels

### v1

@@ Abstract
Recent advances in reinforcement learning (RL) have predominantly leveraged neural network-based policies for decision-making, yet these models often lack interpretability, posing challenges for stakeholder comprehension and trust. Concept bottleneck models offer an interpretable alternative by integrating human-understandable concepts into neural networks. However, a significant limitation in prior work is the assumption that human annotations for these concepts are readily available during training, necessitating continuous real-time input from human annotators. To overcome this limitation, we introduce a novel training scheme that enables RL algorithms to efficiently learn a concept-based policy by only querying humans to label a small set of data, or in the extreme case, without any human labels. Our algorithm, LICORICE, involves three main contributions: interleaving concept learning and RL training, using a concept ensembles to actively select informative data points for labeling, and decorrelating the concept data with a simple strategy. We show how LICORICE reduces manual labeling efforts to to 500 or fewer concept labels in three environments. Finally, we present an initial study to explore how we can use powerful vision-language models to infer concepts from raw visual inputs without explicit labels at minimal cost to performance.
@@ 1 Introduction
In reinforcement learning (RL), agents are tasked with learning a policy, a rule that makes sequential, reactive decisions in complex environments. In recent RL work, agents typically represent the policy as a neural network, as such representations tend to lead to high performance [17]. However, this choice can come at a cost: such policies are challenging for stakeholders to interpret — particularly when the inputs to the network are also complex, such as high-dimensional sensor data. This opacity can become a significant hurdle, especially in applications where understanding the rationale behind decisions is critical, such as healthcare [32] or finance [15]. In such applications, decisions can have significant consequences, so it is essential for stakeholders to fully grasp the reasoning behind actions in order to confidently adopt or collaborate on a policy.
@@
To address interpretability concerns in the supervised learning setting, recent works have integrated human-understandable concepts into the decision-making process through concept bottleneck models [13, 9]. These models incorporate a bottleneck layer whose units correspond to interpretable concepts, ensuring that the final decisions depend on these concepts instead of just on opaque raw inputs. By training the model both to have high accuracy on the target task and to accurately match experts’ concept labels, these models learn a high-level concept-based representation that is simultaneously meaningful to humans and useful for machine learning tasks. As an example, a concept-based explanation for a bird classification task might include a unit that encodes the bird’s wing color.
More recently, these techniques have been applied to RL by incorporating a concept bottleneck in the policy [11, 33], such that the actions taken by the agent are a function of the human-understandable concepts. However, past work assumes that human-provided concept annotations are accessible in the training loop of RL, which presents a significant challenge. In novel domains, we may only have access to data that is not in an interpretable representation, such as RGB or multispectral satellite imagery, while the RL agent needs to know the concepts present in every state and action it encounters during training — even though RL training sets can often be measured in millions or billions of state-action pairs. As a result, the human labelers would need to be actively involved in the RL training process, providing concept labels for unreasonable numbers of observations. This level of involvement is not only detrimental to the human labeler but also risks introducing potential biases and errors in the model training process due to fatigue [16].
In this work, we aim to address this challenge of requiring frequent human interventions during the training of interpretable policies using RL. We propose LICORICE (Label-efficient Interpretable COncept-based ReInforCEment learning), a novel training paradigm consisting of three main contributions. First, LICORICE interleaves concept learning and RL training: it alternately freezes the network layers corresponding to either the concept learning part or the decision-making part. We believe this scheme improves our ability to learn from limited labeled data by reducing interference between the learning tasks. Additionally, concept learning uses data that is more on-policy, which means it is more relevant and useful for learning accurate concepts that directly impact the decision-making process. Second, LICORICE utilizes concept ensembles to actively select the most informative data points for labeling. By focusing on samples that are predicted to provide the most valuable information for model improvement, this technique substantially reduces the number of labels needed to achieve both high performance and high concept accuracy. Third, LICORICE includes a strategy to decorrelate the concept data collected under the current policy. By generating a diverse set of training data, this approach ensures the data remains unbiased and representative of various scenarios. We demonstrate how these changes yield both higher concept accuracy and higher reward while requiring fewer queries on three environments with image input, including an image input version of CartPole and two Minigrid environments.
Given these results, we ask whether we can further reduce the reliance on manual concept labeling by leveraging the potential of vision-language models (VLMs). This capability is important in scenarios where manual concept annotation is impractical. We present an initial exploration of whether VLMs can further reduce the concept annotation burden. We find that VLMs can indeed serve as concept annotators for some but not all of the above environments. In these cases, the resulting policies can achieve $83$ to $93$% of the optimal performance.
@@ 2.1 Reinforcement Learning
In RL, an agent learns to make decisions by interacting with an environment [27]. The environment is commonly modeled as a Markov decision process [21], consisting of the following components: a set of states $\mathcal{S}$, a set of actions $\mathcal{A}$, a state transition function $T:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ that indicates the probability of transitioning from one state to another given an action, a reward function $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ that assigns a reward for each state-action-state transition, and a discount factor $\gamma\in[0,1]$ that determines the present value of future rewards. The agent learns a policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$, which maps states to actions with the aim of maximizing the expected cumulative discounted reward. We evaluate a policy via its value function, which is defined as $V^{\pi}(s)=\mathbb{E}_{\pi}[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}\mid s_{0}=s],\forall s\in\mathcal{S}$. The ultimate aim in RL is to determine the optimal policy, $\pi^{*}$. To do so, the agent iteratively refines its policy based on feedback from the environment.
2.2 Concept Bottleneck Models
@@
Concept-based explanations have emerged as a common paradigm in explainable AI [20]. They explain a model’s decision-making process through human-interpretable attributes and abstractions. Concept bottleneck models [13] are an example of concept-based explanations: they learn a mapping from samples $x\in X$ to labels $y\in Y$ through two functions, $g$ and $f$. The concept encoder function $g:X\rightarrow C$ maps samples from the input space to an intermediate space of human-interpretable concepts. The label predictor function $f:C\rightarrow Y$ maps samples from the concept space to a downstream task space, such as labels for supervised classification. As a result, the prediction $\hat{y}=f(g(x))$ relies on the input $x$ entirely through the bottleneck $\hat{c}=g(x)$.
Training these models requires a dataset of $X\times C\times Y$, in which each sample consists of input features $x$, a ground truth concept vector $c\in\{0,1\}^{k}$, and a task label $y$. The functions $f$ and $g$ are parameterized by neural networks. These models are trained independently, jointly, or sequentially. The independent training method trains $\hat{f}$ and $\hat{g}$ independently. Critically, $\hat{f}$ is trained using the true $c$, but at test time it takes $\hat{g}(x)$ as input. If directly applied to the RL setting, this approach would require continuous access to ground-truth concepts for training $\hat{f}$, which we would like to avoid. In joint training, the model minimizes the weighted sum of the concept losses with the task performance loss. This paradigm can be problematic, as the losses can interfere with each other, and it may require careful tuning to find the right balance between the two losses. In sequential training, the model first learns $\hat{g}$. It then uses the concept predictions $\hat{g}(x)$ to learn $\hat{f}(\hat{g}(x))$. As we will later show, this setup is problematic for RL, as concepts may only emerge after the policy is sufficiently performant.
@@ 3 LICORICE
We now describe LICORICE, our algorithm for interactively querying for concept labels under a limited concept label budget during RL training. In Section 3.1, we describe the architecture of the concept bottleneck policy. We then detail our iterative training process (Section 3.2) and explain our active learning approach for choosing informative training points (Section 3.3). Finally, in Section 3.4, we detail how we use VLMs as a substitute for human concept labeling.
3.1 Concept Bottleneck Policies
We insert a concept bottleneck layer into the policy network, such that $\pi$ maps from states $\mathcal{S}$ to concepts $C$ to actions $\mathcal{A}$, $\pi=f(g(s))$. These concepts serve as intermediaries in the policy network, which subsequently maps the concept vector to a distribution over possible actions. This setup allows the policy to base its decisions on understandable and meaningful features. We describe additional modifications necessary to accommodate the concept bottleneck in Section A.2. As a result, we can use any RL algorithm as long as we include an additional loss function for concept prediction.
Loss Function
We now describe how we use two loss functions for training in our iterative training scheme, described in more detail in Section 3.2. Because we employ an iterative training scheme, we disentangle the two loss functions — the concept prediction loss and the standard RL loss — to prevent interference between them. In this way, the concept prediction loss $L^{C}$ only affects $g$, the part of the model responsible for predicting concepts, and the standard RL loss $L^{\text{RL}}$ only affects $f$. For training $g$, we employ two types of loss functions depending on the nature of the concepts (MSE for continous concepts; cross-entropy loss for categorical concepts). If the problem requires mixed-type concepts, we could discretize the continuous attributes, converting them into categorical forms suitable for classification. This approach ensures that our method accommodates a diverse range of concept types.
3.2 Training Process
We now present LICORICE, our novel algorithm for optimizing concept-based models for RL within a fixed concept query budget. Our iterative training approach extends the sequential bottleneck training paradigm by incorporating multiple phases. Intuitively, this strategy can be advantageous for obtaining accurate on-policy concept estimates. To collect candidate training data for $g$, we employ our current policy for rollouts. To minimize data point correlation and collect a diverse range of data, we use a decorrelation strategy. To select more informative points to query for concept labels, we use a concept ensemble to calculate disagreement.
@@ 3.2 Training Process
17:     Aggregate datasets: $\mathcal{D}_{\text{train}}\leftarrow\mathcal{D}_{\text{train}}\cup\mathcal{D}^{\prime}_{\text{train}}$, $\mathcal{D}_{\text{val}}\leftarrow\mathcal{D}_{\text{val}}\cup\mathcal{D}^{\prime}_{\text{validation}}$
@@
19:     Freeze $g$ and train $f$ using standard RL training to obtain $\pi_{m+1}$
@@
More specifically, LICORICE (Algorithm 1) proceeds in three main stages within each iteration $m\in M$: data collection, concept ensemble training for data selection, and concept bottleneck policy training. First, in the data collection stage (lines 4 to 8), we collect a dataset of unlabeled concept data $\mathcal{U}_{m}$ by rolling out our current policy $\pi_{m}$. Crucially, during this step, we decorrelate the data to ensure the data points are diverse (line 6). Specifically, we set an acceptance probability $p$ for adding the data point to the dataset, as they are generated by the temporally-correlated RL policy. This prepares a diverse dataset for the next stage, with the aim of promoting the training of a more performant concept ensemble for data selection and, eventually, a more accurate concept bottleneck portion of the policy.
@@
The second stage is concept ensemble training for data selection (lines 9 to 16). In this stage, we train the concept ensemble — consisting of $N$ independent concept models — from scratch on the dataset of training points $\mathcal{D}_{\text{train}}$ that has been aggregated over all iterations (line 10). We use this ensemble to calculate the disagreement-based acquisition function, which evaluates whether each candidate in our unlabeled dataset $\mathcal{U}_{m}$ ought to receive a ground-truth concept label (line 11). This function targets samples where prediction disagreement is highest, detailed in Section 3.3. We then query for $B_{m}$ ground-truth concept labels (line 13) to prepare us for the next stage.
We are now prepared for the third stage: concept bottleneck policy training (lines 17 to 19). We aggregate the concept datasets from the second stage with data from previous iterations, and continue training the concept network $g$ on $\mathcal{D}_{\text{train}}$ (line 18). After early stopping dictates that we stop $g$ training, we freeze $g$ and train $f$ using any standard RL algorithm — for example, PPO — to obtain the policy for collecting data in the next iteration (line 19). With that, we are prepared to start the process again from the first stage.
3.3 Active Concept Learning
We propose to leverage an ensemble of concept models to calculate the disagreement-based acquisition function for actively querying for concept labels. To quantify the concept disagreement of a state $U(s)$, we use two different formulations, depending on the concept learning task.
Classification
For concept classification, we use a query-by-committee [24] approach. When the models produce diverse predictions, it indicates that the instance is difficult and more information would be particularly valuable. Conversely, if all models agree, the instance is likely already well understood, and additional labels would be less beneficial. More specifically, we prioritize points with a high proportion of predicted class labels that are not the modal class prediction (also called the variation ratio [3]). This is given by $U(s)=1-\max_{c\in C}\frac{1}{N}\sum_{i=1}^{N}[\tilde{g}_{i}(s)=c]$, where $\tilde{g}_{i}(s)$ is the concept prediction of the $i$-th model on state $s$.
Regression
For concept regression, we observe that we can instead directly use variance as a measure of disagreement. Specifically, the concept disagreement of a state is quantified by the unbiased estimation of variance $U(s)=\sigma^{2}(s)$ of the predictions across the concept models, due to Bessel’s correction, defined as: $U(s)=\sigma^{2}(s)=\frac{1}{N-1}\sum_{i=1}^{N}(\tilde{g}_{i}(s)-\mu(s))^{2}$, where $\mu(s)=\frac{1}{N}\sum_{i=1}^{N}\tilde{g}_{i}(s)$ is the mean prediction of the $N$ concept models.
3.4 Using Vision-Language Models for Concept Labeling
Equipped with our overall algorithm, we now seek to further reduce human annotation burden. To do so, we turn to VLMs due to their remarkable performance in understanding and generating human-like descriptions of visual content [22]. Within the pipeline of LICORICE, we keep all aspects of our algorithm the same but use the VLM as the concept annotator in the training loop. By doing so, we effectively decrease the human annotation effort to zero, albeit with some labeling inaccuracy introduced.
In our experiments, we use GPT-4o [1], a closed-source model that is possibly the most capable vision-language model in the world at the time of writing the paper. During the training loop of LICORICE, we query GPT-4o each time the algorithm requires a concept label (line 13 in Algorithm 1). As we are using a pre-specified concept set, we design prompts with detailed general descriptions of the scene layout and definitions of all concepts, in a similar fashion to giving labeling instructions to real humans. We then prompt the GPT-4o to obtain the generated labels. In environments where continuous concept values are required, we ask GPT-4o to give as accurate estimates as possible; in environments with concepts that are discrete and more intuitive to label, we provide clear instructions of how to read the concept numbers according to the input image. More details about our prompts can be found in Section A.3.

### v2

# LICORICE: Label-Efficient Concept-Based Interpretable Reinforcement Learning

@@ Abstract
Recent advances in reinforcement learning (RL) have predominantly leveraged neural network policies for decision-making, yet these models often lack interpretability, posing challenges for stakeholder comprehension and trust. Concept bottleneck models offer an interpretable alternative by integrating human-understandable concepts into policies. However, prior work assumes that concept annotations are readily available during training. For RL, this requirement poses a significant limitation: it necessitates continuous real-time concept annotation, which either places an impractical burden on human annotators or incurs substantial costs in API queries and inference time when employing automated labeling methods. To overcome this limitation, we introduce a novel training scheme that enables RL agents to efficiently learn a concept-based policy by only querying annotators to label a small set of data. Our algorithm, LICORICE, involves three main contributions: interleaving concept learning and RL training, using an ensemble to actively select informative data points for labeling, and decorrelating the concept data. We show how LICORICE reduces human labeling efforts to 500 or fewer concept labels in three environments, and 5000 or fewer in two more complex environments, all at no cost to performance. We also explore the use of VLMs as automated concept annotators, finding them effective in some cases but imperfect in others. Our work significantly reduces the annotation burden for interpretable RL, making it more practical for real-world applications that necessitate transparency. Our code is released. [^1] https://github.com/cuizhuyefei/LICORICE
@@ 1 Introduction
In reinforcement learning (RL), agents learn a policy, which is a strategy for making sequential decisions in complex environments. Agents typically represent the policy as a neural network, as this representation tends to yield high performance (Mirhoseini et al., 2021). However, this choice can come at a cost: such policies are challenging for stakeholders to interpret — particularly when the network inputs are also complex, such as high-dimensional sensor data. This opacity can pose a significant hurdle, especially in applications where understanding the rationale behind decisions is critical, such as healthcare (Yu et al., 2021) or finance (Liu et al., 2022). In such applications, decisions can have significant consequences, so it is essential for stakeholders to fully grasp the reasoning behind actions to confidently adopt or collaborate on a policy.
@@
To address interpretability concerns in supervised learning, recent works have integrated human-understandable concepts into neural networks through concept bottleneck models (Koh et al., 2020; Espinosa Zarlenga et al., 2022). These models insert a bottleneck layer whose units correspond to interpretable concepts, ensuring that the final decisions depend on these concepts instead of on opaque raw inputs. By training the model both to have high task accuracy and to accurately match experts’ concept labels, these models learn a high-level concept-based representation that is simultaneously meaningful to humans and useful for machine learning tasks. As an example, a concept-based explanation for a bird classification task might include a unit that encodes the bird’s wing color. Concepts also offer a path to compositional generalization by enabling subtask decomposition into meaningful, reusable components (Mao et al., 2022; Wang et al., 2023).
More recently, these techniques have been applied to RL by incorporating a concept bottleneck in the policy (Grupen et al., 2022; Zabounidis et al., 2023), so that the chosen actions are a function of the human-understandable concepts. However, a significant challenge emerges in this method’s practical application: past work assumes that concept annotations are readily available during RL training. To learn a mapping from states to concepts, an RL agent requires concept information for every state encountered during training, which is often millions or billions of state-action pairs. Because many domains do not automatically provide these concept labels (e.g., autonomous driving), we require a means of obtaining them. Relying on human labelers is impractical, risking errors due to fatigue (Marshall & Shipman, 2013). Using vision language models (VLMs) for automated concept extraction (Oikarinen et al., 2023), while alleviating the human bottleneck, introduces significant API or inference costs that can be prohibitive.
Figure 1: LICORICE overview. In concept-based RL, the policy first maps from states to the concepts in a bottleneck layer with $g$, and then maps from concepts to (distributions over) actions with $f$. During training, LICORICE addresses concept label efficiency concerns with three key components: i) iterative training, ii) data decorrelation, and iii) active learning.
In this work, we address this challenge with LICORICE (Label-efficient Interpretable COncept-based ReInforCEment learning), a novel training paradigm designed to minimize the number of concept annotation queries while maintaining high task performance. Figure 1 illustrates our algorithm. LICORICE tackles three key challenges. First, it addresses the problem of concept learning on off-policy or outdated data. If concepts are learned from data collected by a random policy, the concept distribution may not reflect the data from an optimal policy. To ensure that concept learning occurs on more recent and on-policy data, LICORICE interleaves concept learning and RL training through iterative training: it alternately freezes the network layers corresponding to either the concept learning part or the decision-making part. Second, LICORICE addresses the problem of limited training data diversity that occurs when an agent interacts with the environment, thereby generating sequences of highly similar, temporally correlated data points. To tackle this issue, LICORICE implements a data decorrelation strategy to produce a more diverse set of training samples. Third, LICORICE addresses the inefficient use of annotation effort, where labeled samples may provide redundant information. To resolve this problem, we employ disagreement-based active learning using a concept ensemble to select the most informative data points for labeling.
To evaluate the effectiveness of LICORICE, we conduct experiments in two scenarios—perfect human annotation and VLM annotation—on five environments with image input: an image-based version of CartPole, two Minigrid environments, and two Atari environments. First, with assumption of perfect human annotation, we show that LICORICE yields both higher concept accuracy and higher reward while requiring fewer annotation queries compared to baseline methods. Second, we find that VLMs can indeed serve as concept annotators for some, but not all, of the environments.
Our contributions are summarized as follows:
  • To the best of our knowledge, we are the first to investigate the problem of a limited concept annotation budget for interpretable RL. We introduce LICORICE, a novel training scheme that enables label efficient learning of concept-based RL policies.
  • We conduct extensive experiments to show the effectiveness of LICORICE across five environments with varying budget constraints.
  • We study the use of VLMs as automated concept annotators, demonstrating their effectiveness in certain environments while highlighting challenges in others.
@@ 2 Preliminaries
Reinforcement Learning. In RL, an agent learns to make decisions by interacting with an environment (Sutton & Barto, 2018). The environment is commonly modeled as a Markov decision process (Puterman, 2014), consisting of the following components: a set of states $\mathcal{S}$, a set of actions $\mathcal{A}$, a state transition function $T:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ that indicates the probability of transitioning from one state to another given an action, a reward function $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ that assigns a reward for each state-action-state transition, and a discount factor $\gamma\in[0,1]$ that determines the present value of future rewards. The agent learns a policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$, which maps states to distributions over actions with the aim of maximizing the expected cumulative discounted reward. We evaluate a policy via its value function, which is defined as $V^{\pi}(s)=\mathbb{E}_{\pi}[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}\mid s_{t}=s]$. The ultimate aim in RL is to determine the optimal policy, $\pi^{*}$, through iterative refinement based on environmental feedback.
@@
Concept Policy Models. Concept-based explanations have emerged as a common paradigm in explainable AI (Poeta et al., 2023). They explain a model’s decision-making process through human-understandable attributes and abstractions. In supervised learning, concept bottleneck models (Koh et al., 2020) implement this approach using two key functions: a concept predictor $g:X\rightarrow C$, mapping inputs to interpretable concepts, and an output predictor $f:C\rightarrow Y$, mapping the concept predictions to a downstream task space, such as labels for classification. As a result, the prediction takes the form $\hat{y}=f(g(x))$, so that the input $x$ influences the output solely through the bottleneck layer $\hat{c}=g(x)$. In RL, since we want the policy to be interpretable, we insert a concept bottleneck layer into the policy network:
$\pi(s)=f(g(s)),$
such that $\pi$ maps from states $s\in\mathcal{S}$ to concepts $c\in C$ to actions $a\in\mathcal{A}$ (Grupen et al., 2022; Zabounidis et al., 2023). This setup allows the policy to base its decisions on understandable and meaningful features. As a result, we can use any RL algorithm that can be modified to include an additional loss function for concept prediction.
Training Concept Policy Models. Training these models requires a dataset of state-concept-action triplets $(s,c,a)\in\mathcal{S}\times C\times\mathcal{A}$. The functions $f$ and $g$ are typically implemented as neural networks, with their parameters collectively defined as $\theta$. Previous work (Zabounidis et al., 2023) simply combines the concept prediction loss $L^{C}(\theta)$ and RL loss $L^{\text{RL}}(\theta)$:
$L(\theta)=L^{\text{RL}}(\theta)+\lambda_{C}L^{C}(\theta),$
where the exact definitions of $L^{\text{RL}}(\theta)$ and $L^{C}(\theta)$ depend on the choice of RL algorithm and concept learning task. The objective is to find the optimal parameters $\theta^{*}$ that minimize this combined loss function with the coefficient $\lambda_{C}$. However, this approach requires continuous access to ground-truth concepts for training $f$, which may not always be feasible or desirable in practical RL scenarios.
@@ 3 LICORICE
As we have mentioned, the standard way of training concept-based RL assumes continuous access to an oracle to provide concept labels. However, this assumption is problematic due to the large annotation cost incurred by human or automated labeling efforts. To reduce the number of concept labels required for concept-based RL, we propose LICORICE, a novel algorithm for interpretable RL consisting of three main components: iterative training, data decorrelation, and disagreement-based active learning. The full pseudocode is in Algorithm 1.
Iterative Training. As the agent’s policy improves, the distribution of the visited states and their associated concepts changes. Consider an MDP where states are indexed. With a random (initial) policy, the agent tends to visit small-index states near the initial state, encountering only the concepts relevant to those states. However, as the policy improves, the agent explores higher-index states, leading to a shift in both the state and concept visitation distribution. If we exhaust all queries at the beginning of training, we risk training the model only on the concepts associated with small-index states from the random policy, potentially missing important concepts that emerge as the agent explores. We therefore propose iterative training to enable LICORICE to progressively refine its understanding of concepts as the policy improves. Iterative training consists of two parts: concept learning and behavior learning. Assume that we have access to an annotated concept dataset $\mathcal{D}_{\text{train}}$ generated by our current policy $\pi_{m}$. Then, concept learning focuses on training the concept portion of the network $g$ on this dataset $\mathcal{D}_{\text{train}}$ (line 18). Behavior learning involves freezing $g$ and training the behavior portion of the network $f$ with any standard RL algorithm with its associated loss $L^{\text{RL}}$ (line 19). Upon completion of behavior learning, we obtain an updated policy $\pi_{m+1}$, which serves to collect more unlabeled concept data to help train $g$ in the next iteration.
@@
17:     Aggregate datasets: $\mathcal{D}_{\text{train}}\leftarrow\mathcal{D}_{\text{train}}\cup\mathcal{D}^{\prime}_{\text{train}}$, $\mathcal{D}_{\text{val}}\leftarrow\mathcal{D}_{\text{val}}\cup\mathcal{D}^{\prime}_{\text{val}}$
@@
19:     Freeze $g$ and continue training $f$ using standard RL training to obtain $\pi_{m+1}$
@@
Data Decorrelation. We do not obtain ground-truth concept labels for all states when rolling out the current policy $\pi_{m}$. Instead, rollouts produce a dataset $\mathcal{U}_{m}$ of unlabeled states as candidates for querying for ground-truth concepts from an oracle. Simply collecting and labeling all encountered states would give us long chains of nearly-identical samples, undermining the diversity we need for effective learning. To resolve this challenge, we introduce data decorrelation with two key components: a budget multiplier $\tau$ that sets $|\mathcal{U}_{m}|=\tau\cdot B_{m}$, and a per-state acceptance probability $p$. This random acceptance/rejection mechanism leverages a fundamental property of Markov processes—states become increasingly independent as their temporal distance grows according to the mixing time—allowing us to build more diverse datasets from our policy rollouts.
@@
Disagreement-Based Active Learning. Equipped with $\mathcal{U}_{m}$, we are now prepared to select data points for querying for concept labels. The purpose of this stage is to make use of our limited labeling budget $B_{m}$ to collect a labeled dataset $\mathcal{D}_{m}$ for training $g$. We propose to train a concept ensemble—consisting of $N$ independent concept models—from scratch on the dataset of training points $\mathcal{D}_{\text{train}}$ that has been aggregated over all iterations (line 10). We use this ensemble to calculate the disagreement-based acquisition function $\alpha(\cdot)$, which determines whether each candidate in $\mathcal{U}_{m}$ ought to receive a ground-truth concept label (line 11). This function targets samples where prediction disagreement is highest among ensemble members, as these points often represent areas of uncertainty or decision boundaries where additional labeled data would be most informative. For $\alpha(\cdot)$, we use two different formulations, depending on the concept learning task. For concept classification, we use a query-by-committee approach (Seung et al., 1992), in which we prioritize points with a high proportion of predicted class labels that are not the modal class prediction (also called the variation ratio (Beluch et al., 2018)):
$\alpha(s)=1-\max_{c\in C}\frac{1}{N}\sum_{i=1}^{N}[\tilde{g}_{i}(s)=c].$
Here, $\tilde{g}_{i}(s)$ is the concept prediction of the $i$-th model on state $s$. For concept regression, we directly use the estimate of variance across the concept models as a measure of disagreement:
$\alpha(s)=\sigma^{2}(s)=\frac{1}{N-1}\sum_{i=1}^{N}(\tilde{g}_{i}(s)-\mu(s))^{2},$
where $\mu(s)=\frac{1}{N}\sum_{i=1}^{N}\tilde{g}_{i}(s)$ is the mean prediction of the $N$ concept models. After querying for $B_{m}$ ground-truth concept labels (line 13) using $\alpha(\cdot)$, we are prepared to continue training $g$ during the concept learning stage.
Implementation Details. For training $g$ (line 18), we employ two types of loss functions depending on the nature of the concepts: MSE for continuous concepts and cross-entropy loss for categorical concepts. If the problem requires mixed-type concepts, we either discretize continuous attributes or simply use a mixed loss. To train $f$ (line 19), we freeze $g$ and use PPO to continue training $f$ from the previous iterations, resulting in the final policy $\pi_{M+1}$. For complex environments that require many iterations, we observe that RL-related parameters may get stuck in a local region in early iterations, becoming hard to optimize later. To mitigate this issue, in the last iteration, we additionally train a new RL network $f^{\prime}$ from scratch when fixing $g$ with $\pi_{M+1}$ as the anchoring policy to get the final $\pi^{\prime}_{M+1}$. The anchoring policy ensures $\pi^{\prime}_{M+1}$ has a similar distribution to the previous one and the annotated observations are still highly relevant. Specifically, we initialize another policy $\pi_{\theta}$ with the same frozen $g$ parameters and randomly initialized $f$, then train it with the standard PPO loss function and an additional KL-divergence penalty between the current policy and $\pi_{M+1}$:
$L(\theta)=L^{\text{PPO}}(\theta)+\beta\cdot\text{KL}[\pi_{M+1}(\cdot\mid s),\pi_{\theta}(\cdot\mid s)].$